A stone of mass 50kg released from a heigth of 2m above the ground level if the stone falls freely to a heigth of 5m above the ground, its velocity is?
1
Expert's answer
2019-06-03T09:55:20-0400

From the conservation of enegy:


"mgH=mgh+0.5mv^2"

The velocity:


"v=\sqrt{2g(H-h)}"

"v=\sqrt{2(9.8)(20-5)}=17\frac{m}{s}"